(1) Brown's gross pay is $127.50 per week. He works 30 hours a week. How much does Brown earn per hour?
(2) one week, Brown has only three deductions: Federal withholding tax of $15.94, State withholding tax of $5.10, amd F.I.C.A of $8.03.How much do these three deductions total?
(3)After these three items are deducted, what will Brown's net pay be?
(4) The employees where Brown works have a new contract for an 11% raise. What will be his gross pay after 11% raise? ( answer requires rounding to the hearest cent)
(5) Brown must pay a 6% F.I.C.A. Tax on his new gross pay. How much will be left in his pay after this F.I.C.A is deducted? ( use answer #4 above).

Answer by checkley71(8403)

1) 124.25/30=$4.25 per hour
2) 15.94+5.10+8.03=$29.07
3) 124/25-29.07=$95.18
4) 4.25*1.11=$4.72 per hour
5) 4.72*30=141.60 minus .06*141.60=$8.50 or his pay after deducting his new FICA is 141.60-8.50=$133.10